# Suppression of errors in collectively coded information

### v1

@@ II.2 Mutant suppression in the simplified model
Following simulation of repeated thermal melt-anneal-extend cycles, we first observe that oligos extend on each other (Fig. 4B, top); the distribution of oligo lengths shifts from their initial values (25 bp) to a maximal length set by a combination of $o_{max}$ (the maximal duplex overlap length for unbinding during the melt phase) and $l_{max}$ (the maximal oligo length). As a consequence of their extension, oligos become longer and more likely to be trapped in duplexes that overlap too much to unbind during the melting step. In particular, while oligos at the end of thermal cycle 0 are still all single-stranded, virtually all oligos are bound in duplexes of length 55-60 bp at the end of cycle 10 (Fig. 4B, bottom). We refer to this state where all oligos are duplex-bound as ‘pool stasis’, since the VCG cannot replicate anymore. We note that introducing a source of short oligos (e.g. from external sources or from templated nucleation of activated nucleotides) would instead allow for a continued steady state of pool replication.
@@
To model the qPCR readouts of the DNA model experiments, we track the relative amplification of the mut and wt alleles by measuring the total concentration of oligos that contain the allele along with a flanking region around the allele. This flanking region is set by the oligo that the mut allele initially appears on. After normalizing these concentrations to the initial concentration of flanking region oligos, we can define a ‘wt advantage’ statistic as the ratio between the wt region amplification at pool stasis and the mut region amplification at pool stasis. We find that the wt advantage increases with the virtualness $V$ of the initial oligo pool (Fig. 4C), where $V$ here can be defined as the number of unique oligos that cover the consensus VCG sequence in the oligo pool. We also find that the wt advantage measured at pool stasis is stronger as the initial mutant position shifts from the 5 end to the 3 of the oligo it starts on (Fig. 4D).
@@ II.3 Collective binding partner effects predict mutant suppression
Figure 5: A binding partner effect explains the suppression of initially 3-proximal mutants. (A) A mutant allele located towards the 3 end of an oligo has only 1 ‘productive’ binding partner capable of copying that mutant allele through extension (upper schematic) while an allele located at the 5 end has 3 productive binding partners capable copying it (lower schematic). Hence 5-proximal alleles are copied more rapidly. (B) Evolution of distributions for mut $P^{mut}(x)$ and wt $P^{wt}(x)$ allele distance $x$ from 3 end of oligos that contain the allele over thermal cycles. Distribution is for single-stranded oligos only. Relative concentration across each column is normalized by max binned concentration value. (C) (left) Average distance of mut and wt allele from 3 end of oligos that contain allele as a function of thermal cycle. (right) Evolution of the average productive pairing fraction $f^{A}$ for mutant and wt alleles as a function of thermal cycle. For oligo $i$ containing allele AA, we define a productivity factor $F^{A}_{i}$ as the ratio between the concentration of single-stranded oligos that can bind to oligo $i$ downstream of $A$ and the total concentration of oligos which can bind to oligo $i$. For oligos only bound in duplexes, $F^{A}_{i}=0$. The productive pairing fraction $f^{A}$ is a concentration-weighted average over $F^{A}_{i}$.
@@
In order to find an explanation for the observation of mutant suppression both in our experimental DNA system as well as our simplified simulation model, we focus on the case where the mutant allele is initially at the 3 end of the oligo on which it occurs. This is both the situation where the mutant suppression is the most dramatic (Fig. 4D), as well as the most relevant for mutant alleles that arise through mutations during extension.
@@
We propose that the wt advantage comes from a binding partner effect. In the VCG pool, each oligonucleotide has multiple potential reverse complement binding partners, with the number of potential partners increasing with the virtualness $V=3,\ldots 12$ of the initial VCG oligo pool. These binding partners all compete for annealing with one another on the template oligo, on equal footing with the perfect reverse complement partner.
@@
Crucially, only binding partners whose 3 ends anneal downstream of the allele (i.e., between the allele and the 3 end of the template), can increase that allele’s reverse complement concentration via extension. In contrast, binding partners whose 3 ends are upstream of the allele may extend but do not result in copying of the allele’s reverse complement, precluding an increase in concentration of the allele itself upon further rounds of extension.
@@
Therefore, a mutant allele located near the 3 end of an oligo has fewer binding partners that can bind downstream of that allele and make its reverse complement during extension. In comparison, the same allele, when located towards the 5 end of an oligo, can be copied off of by almost all binding partners that anneal to that template oligo (see Fig. 5A).
@@
The binding partner effect implies that the instantaneous per-cycle rate at which an allele on a given oligo can produce more copies of its reverse complement is set by two quantities: 1. the concentration of ‘productive’ binding partner oligos which can bind and extend past the allele and are thus productive in copying that allele, and; 2. the concentration of all binding partner oligos that compete with (1), regardless of whether they can pick up the allele upon extension. An allele is therefore more productive if it exists further from the 3 oligo end and hence a greater fraction of the oligo’s binding partners can form pairings that allow the allele to be copied. Correspondingly, an allele is also more productive if it sits closer to the 5 oligo end since there are fewer binding partners that do not pick up the allele upon extension.
@@
where $c_{ss,j}$ is the concentration of single-stranded oligo $j$, $D^{A}_{i}$ is the set of oligos that bind with 3-ends downstream of the allele $A$ on oligo $i$, and $B_{i}$ is the set of all oligos that can bind to oligo $i$ regardless of the configuration of the duplex they form. (When oligo $i$ only exists in duplexes, $F^{A}_{i}$ is defined to be 0.)
@@
This definition captures the intuition of how the position of an allele on an oligo influences the allele’s productivity. For instance, when oligos are shorter than the oligo length maximum $l_{max}$, an allele $A$ coded on the 5 end of oligo $i$ will have a maximal productivity factor $F^{A}_{i}=1$. This high productivity comes because every binding partner of oligo $i$ is capable of copying $A$ through extension, so every binding partner oligo index $j\in D^{A}_{i}$ is also $j\in B_{i}$. In contrast, $F^{A}_{i}=0$ for that same allele $A$ if it were coded at the 3 end of oligo $i$; in this case, there are no downstream binding partners for $A$ on $i$, so $D^{A}_{i}$ is empty. An approximation of the link between allele position and productivity that interpolates between the 3- and 5-end location limits is therefore:
@@
where $x^{A}_{i}$ is the distance of allele $A$ from the 3 end of oligo $i$, and $l_{i}$ is the length of oligo $i$.
Therefore, in order to understand the time evolution of $f^{A}$ for the wt and mut alleles, we should track the distributions $P^{mut}(x)$ and $P^{wt}(x)$ of an allele’s distance $x$ from the 3 end of oligos across single-stranded oligos in the oligo pool (Fig. 5B). While the mutant allele is initially 3-proximal (concentrated at 5 bp from the 3 end), by cycle 10 $P^{mut}(x)$ eventually spreads out. $P_{wt}(x)$, in contrast, is by construction initially more spread out since the wt allele was already encoded on a variety of different oligos (e.g. see Fig. 2A), and remains uniformly spread after all 10 cycles. Tracking the average distance $\langle x\rangle$ of the mut and wt alleles from the 3 end (Fig. 5C, left) reveals that the mutant allele is initially closer to the 3 end of oligos compared to the average wt allele, indicating an initial disadvantage in replicative potential. However, as the oligo pool replicates, the relative positions of the mean wt and mean mutant allele converge.
Indeed, if we calculate the time-dependent productive pairing fraction $f^{A}$ (Fig. 5C, right), we find that mut alleles are initially at a disadvantage, with a lower productive pairing fraction $f^{mut}$ compared to wt allele $f^{wt}$. This disadvantage decreases over the first few replication cycles, consistent with the decreasing gap between mean mut and wt distances from oligo 3 ends. By cycle 5, the productive pairing fraction $f^{A}$ of wt and mutant alleles are essentially equal and both very close to 0. The collapse of the productive pairing fraction $f^{A}$ for both mut and wt alleles is due to the fact that most alleles are bound in fully-extended blunt end duplexes. Such duplexes cannot participate in subsequent rounds of melt-anneal-extend and hence cannot contribute to the production of more mut or wt alleles.
In summary, the initial disadvantage faced by 3-proximal mutant alleles due to the binding partner effect is transient. However, the cumulative impact of this early disadvantage leads to a persistent suppression of the mutant allele relative to the wildtype, even by the time of pool stasis. We next discuss how the stalling effect[19, 20, 15, 21, 16, 22] can significantly prolong this transient phase, thereby amplifying the overall suppression of the mutant allele.
@@ III Stalling enhances collective suppression of mutants
Figure 6: Combination of collective effects and stalling suppresses mutant growth. (A) In duplexes where priming oligo has a correctly matched 3 end, extension proceeds normally and all such primers fully extend to form blunt ends. In duplexes where priming oligo has an incorrectly matched 3 end, extension past the mismatch takes additional time, $\tau_{stall}$. Consequently, only a fraction $1-e^{-\frac{\tau_{cycle}}{\tau_{stall}}}$ of mismatched oligos manage to extend within a thermal cycle of time $\tau_{cycle}$. See Supp. Sec. VIII for further details. (B) (top) mut and wt amplification for three different values of $\frac{\tau_{stall}}{\tau_{cycle}}$ as a function of thermal cycle. (bottom) Mutant allele amplification measured after 10 thermal cycles as a function of $\frac{\tau_{stall}}{\tau_{cycle}}$.
The stalling effect[2] describes the reduced rate of extension from a primer that ends in a mismatched base pair, compared to extension from a perfectly matched primer. Stalling has been observed across a wide range of templated replication systems, including complex DNA and RNA polymerases[21, 23, 24], ribozymes[25], and even non-enzymatic RNA extension[16]. It has been proposed as a general mechanism for mutant suppression[16, 26] and as a potential evolutionary precursor to more elaborate error-correcting proofreading systems[22]. Here, we investigate how stalling influences the propagation of mutant alleles within the VCG framework.
We consider annealed configurations where one of the oligos in the duplex is not correctly matched to its template at its 3 end. This can either happen when an oligo with a mutant allele at its 3 end is annealed to an oligo with a wt allele at that position, or it is an oligo with a wt allele at its 3 end annealed to an oligo with a mutant allele there.
For such duplexes where one of the oligos has a mismatch at its 3 end, stalling delays extension by a time $\tau_{stall}$. Whether or not the duplex gets to extend therefore depends crucially on whether thermal cycling (with timescale $\tau_{cycle}$) drives the duplex apart before the stalling delay can be overcome. For primers with no 3 mismatch, there is no barrier to extension and such primers will fully extend during the current thermal cycle. To model the effect of stalling, we posit that primers with a mismatched 3 end get to extend through an exponential decay process with timescale $\tau_{stall}$. Hence by the end of a melt-anneal-extend cycle of length $\tau_{cycle}$, only a fraction $1-e^{-\frac{\tau_{cycle}}{\tau_{stall}}}$ of primers with a mismatched 3 end will have extended (Fig. 6A). Please see Supp Sec. VIII for further discussion and implementation details on stalled extension.
@@ IV Discussion
Our results show that collective replication in a virtual circular genome (VCG) architecture intrinsically suppresses the propagation of mutant alleles, even in the absence of high-fidelity copying or kinetic proofreading. Suppression is strongest when mutations are 3-proximal and increases with the VCG’s ‘virtualness’ (redundancy of overlapping oligos). Experiments, simulations, and a reduced theory agree that mutant alleles are selectively under-amplified relative to wildtype alleles due to competitive binding for productive extension partners.
@@
Mechanistically, the suppression follows from a binding partner effect: an allele’s chance of being copied depends on how many annealed partners place their 3 end downstream of that allele so that extension traverses it. 3-proximal mutants have fewer such productive partners and thus are at a disadvantage compared to wildtype alleles that are coded at varying distances from the 3 end. Polymerase stalling at mismatched 3 ends prolongs this period of lowered replicative ability by delaying extension past the mismatch. Consequently, the combined effect of partner limitation and stalling leads to strong suppression that strengthens with virtualness and with increasing ratio $\tau_{stall}/\tau_{cycle}$ of stalling time to thermal melt-anneal-extend cycle time. Unlike mispriming suppression in PCR, which relies on reduced binding affinity from mismatches, our effect persists even when binding between wildtype and mutant is allowed.
@@
Prior classic theories of distributed replication (e.g., hypercycles[1]) posit cooperative networks that can stabilize information, but they have been difficult to realize and manipulate at the molecular level[27, 28]. Although the VCG can perhaps be cast, in retrospect, as a cooperative network in the spirit of hypercycle theory, our work begins from a concrete molecular specification: here, a set of overlapping strands and their allowed anneal–extend interactions. That molecular starting point, rather than an abstract catalytic graph, is what makes the system experimentally tractable and tunable, allowing controlled variation of virtualness, mutation position, and kinetic parameters. In this aspect, our system shares features with molecularly-defined experimental models[29, 27, 30, 31, 32]. While the focus of these prior works was often the emergence of catalytic cycles from random sequence pools, we focused on the maintenance of information in a more structured pool. This assumed structure let us quantify an architecture-driven fidelity mechanism and map how virtualness and 3 proximity govern the fate of mutants.
Our experiments emphasize primer extension in a fixed pool of oligomers that eventually ‘die’ by forming long, unmeltable duplexes; sustained replication will require continual regeneration of short oligos (e.g., de novo synthesis from activated nucleotides). Moving to driven steady states could strengthen mutant suppression, especially under strong stalling, because wildtype alleles can continue to amplify while 3-proximal mutants remain suppressed. While we focused on templated extension, templated ligation is natural in distributed architectures [8, 9, 10, 33, 34] and enables growth from both 5 and 3 ends. The binding partner effect will have to be studied and quantified in the context of ligation to understand biases intrinsic to that mode of replication. We expect the core ecological mechanism revealed by our experimental DNA model to carry over to RNA; however, backbone-specific chemistry (persistence length, hydrolysis, non-enzymatic extension kinetics) will shift quantitative outcomes. Additional instabilities, most notably mismatched annealing[35] and sequence repeats[36], remain to be mapped to determine parameter regimes in which VCGs are stable. In particular, some mutations could create repeats which will allow oligomers to anneal and extend off novel regions of the genome, generating novel chimeric oligomers. Finally, our complete set of results on mutations at varying distances from the end also explain the fate of mutations that arise due to processes like deamination[37] which are not confined to the 3 end.
@@
Despite these limitations and potential extensions, our work shows that genomic architecture can act as an intrinsic bias for how stored information evolves. That is, by filtering which novelties are propagated versus rejected, the architecture shapes evolvability and effectively tunes the error threshold without explicit proofreading. Comparable architecture-level filters might be operating in extant biology, for example how RNA viruses with segmented genomes can undergo genetic reassortment[38]. Other non-standard genome architectures include those of multipartite bacteria (e.g., Borrelia burgdorferi[39, 40]), which partition essential genes across a chromosome plus several plasmids; and trypanosomatids[41, 42], which maintain kinetoplast DNA as a topologically linked network of maxi- and minicircles. In all these cases, information is distributed across many molecules with heterogeneous maintenance and inheritance, rather than on one continuous polymer. We lack a general framework for how such architectures influence mutation fate, dosage control, recombination, assortment noise, and long-term evolvability or why particular architectures arise and persist. While we do not address architectural features specific to these extant systems, our work highlights open questions about the trade-offs of distributing genomic information across many versus few molecules.

### v2

@@ II.2 Mutant suppression in the simplified model
Following simulation of repeated thermal melt-anneal-extend cycles, we first observe that oligos extend on each other (Fig. 4B, top); the distribution of oligo lengths shifts from their initial values (25 nt) to a maximal length set by a combination of $o_{max}$ (the maximal duplex overlap length for unbinding during the melt phase) and $l_{max}$ (the maximal oligo length). As a consequence of their extension, oligos become longer and more likely to be trapped in duplexes that overlap too much to unbind during the melting step. In particular, while oligos at the end of thermal cycle 0 are still all single-stranded, virtually all oligos are bound in duplexes of length 55–60 bp at the end of cycle 10 (Fig. 4B, bottom). We refer to this state where all oligos are duplex-bound as ‘pool stasis’, since the VCG cannot replicate anymore. In simulations, we see pool stasis occurs at approximately cycle 10.
@@
To model the qPCR readouts of the DNA model experiments, we track the relative amplification of the mut and wt alleles by measuring the total concentration of oligos that contain the allele along with a flanking region around the allele. This flanking region is set by the oligo that the mut allele initially appears on. After normalizing these concentrations to the initial concentration of flanking region oligos, we can define a ‘wt advantage’ statistic as the ratio between the wt region amplification at pool stasis and the mut region amplification at pool stasis. We find that the wt advantage increases with the virtualness $V$ of the initial oligo pool (Fig. 4C), where $V$ here can be defined as the number of unique oligo pairs that cover the consensus VCG sequence in the oligo pool (either 3, 6, or 12). We also find that the wt advantage measured at pool stasis is stronger as the initial mutant position shifts from the 5 end to the 3 of the oligo it starts on (Fig. 4D).
@@ II.3 Collective binding partner effects predict mutant suppression
Figure 5: A binding partner effect explains the suppression of initially 3-proximal mutants. (A) A mutant allele located towards the 3 end of an oligo has only 1 ‘productive’ binding partner capable of copying that mutant allele through extension (left schematic). In contrast, the wt allele is present at varying positions on different oligos, including 5 end placements that have many productive binding partners capable of copying it (right schematic). (B) Evolution of distributions for mut $P^{mut}(x)$ and wt $P^{wt}(x)$ allele distance $x$ from 3 end of oligos that contain the allele over thermal cycles. Distribution is for single-stranded oligos only. Relative concentration across each column is normalized by max binned concentration value. (C) (left) Average distance of mut and wt allele from 3 end of oligos that contain allele as a function of thermal cycle. (right) Evolution of the average productive pairing fraction $f^{A}$ for mutant and wt alleles as a function of thermal cycle. For oligo $i$ containing allele AA, we define a productivity factor $F^{A}_{i}$ as the ratio between the concentration of single-stranded oligos that can bind to oligo $i$ downstream of $A$ and the total concentration of oligos which can bind to oligo ii. For oligos only bound in duplexes, $F^{A}_{i}=0$. The productive pairing fraction $f^{A}$ is a concentration-weighted average over $F^{A}_{i}$.
@@
In order to find an explanation for the observation of mutant suppression both in our experimental DNA system as well as our simplified simulation model, we focus on the case where the mutant allele is initially at the 3 end of the oligo on which it occurs. This is both the situation where the mutant suppression is the most dramatic (Fig. 4D), as well as the most relevant for mutant alleles that arise through mistakes during extension.
@@
We propose that the wt advantage comes from a binding partner effect. In the VCG pool, each oligomer has multiple potential reverse complement binding partners, with the number of potential partners increasing with the virtualness $V=3,\ldots,12$ of the initial VCG oligo pool. These binding partners all compete for annealing with one another on the template oligo, on equal footing with the perfect reverse complement partner.
@@
Crucially, only binding partners whose 3 ends anneal downstream of the allele (i.e., between the allele and the 3 end of the template), can increase that allele’s reverse complement concentration via extension. In contrast, binding partners whose 3 ends are upstream of the allele may extend but do not result in copying of the allele into its reverse complement, precluding an increase in concentration of the allele itself upon further rounds of extension.
@@
Therefore, a mut allele initially located near the 3 end of an oligo has fewer binding partners that can bind downstream of that allele and make its reverse complement during extension. In contrast, wt alleles are initially distributed across many different oligo locations. When the wt allele is located towards the 5 end of a template oligo, it can be copied by almost all binding partners that anneal to that template (see Fig. 5A).
@@
The binding partner effect implies that the instantaneous per-cycle rate at which an allele on a given oligo can produce more copies of its reverse complement is set by two quantities: 1. the concentration of ‘productive’ binding partner oligos which can bind and extend past the allele and are thus productive in copying that allele, and; 2. the concentration of all binding partner oligos that compete with (1), regardless of whether they can pick up the allele upon extension. An allele is more productive if it exists further from the 3 oligo end and hence more of the oligo’s binding partners can form pairings that allow the allele to be copied. Correspondingly, an allele is also more productive if it sits closer to the 5 oligo end since there are fewer binding partners that do not pick up the allele upon extension.
@@
where $c_{ss,j}$ is the concentration of single-stranded oligo $j$, $D^{A}_{i}$ is the set of oligos that bind with 3 ends downstream of the allele $A$ on oligo $i$, and $B_{i}$ is the set of all oligos that can bind to oligo $i$ regardless of the configuration of the duplex they form. (When oligo $i$ only exists in duplexes, $F^{A}_{i}$ is defined to be 0.)
@@
This definition captures the intuition of how the position of an allele on an oligo influences the allele’s productivity. For instance, when oligos are shorter than the oligo length maximum $l_{max}$, an allele $A$ coded on the 5 end of oligo $i$ will have a maximal productivity factor $F^{A}_{i}=1$. This high productivity comes because every binding partner of oligo $i$ is capable of copying $A$ through extension, so every binding partner oligo index $j\in D^{A}_{i}$ is also $j\in B_{i}$. In contrast, $F^{A}_{i}=0$ for that same allele $A$ if it were coded at the 3 end of oligo $i$; in this case, there are no downstream binding partners for $A$ on $i$, so $D^{A}_{i}$ is empty. An approximation of the link between allele position and productivity that interpolates between the 3 and 5 end location limits is therefore:
@@
where $x^{A}_{i}$ is the distance of allele $A$ from the 3 end of oligo $i$, and $l_{i}$ is the length of oligo $i$; more generally, $F^{A}_{i}$ is a monotonically increasing function of $x^{A}_{i}$ which is 0 at $x^{A}_{i}=0$.
Therefore, in order to understand the time evolution of $f^{A}$ for the wt and mut alleles, we should track the distributions $P^{mut}(x)$ and $P^{wt}(x)$ of an allele’s distance $x$ from the 3 end of oligos across single-stranded oligos in the oligo pool (Fig. 5B). While the mutant allele is initially 3-proximal (concentrated at 5 nt from the 3 end), by cycle 10 $P^{mut}(x)$ eventually spreads out. $P_{wt}(x)$, in contrast, is by construction initially more spread out since the wt allele was already encoded on a variety of different oligos (e.g. see Fig. 2A), and remains uniformly spread after all 10 cycles. Tracking the average distance $\langle x\rangle$ of the mut and wt alleles from the 3 end (Fig. 5C, left) reveals that the mutant allele is initially closer to the 3 end of oligos compared to the average wt allele, indicating an initial disadvantage in replicative potential. However, as the oligo pool replicates, the relative positions of the mean wt and mean mutant allele converge.
Indeed, if we calculate the time-dependent productive pairing fraction $f^{A}$ (Fig. 5C, right), we find that mut alleles are initially at a disadvantage, with a lower productive pairing fraction $f^{mut}$ compared to wt allele $f^{wt}$. This disadvantage decreases over the first few replication cycles, consistent with the decreasing gap between mean wt and mut distances from oligo 3 ends. By cycle 5, the productive pairing fraction $f^{A}$ of wt and mut alleles are essentially equal and both very close to 0. The collapse of the productive pairing fraction $f^{A}$ for both wt and mut alleles is due to the fact that most alleles are bound in fully-extended blunt end duplexes. Such duplexes cannot participate in subsequent rounds of melt-anneal-extend and hence cannot contribute to the production of either wt or mut alleles.
In summary, the initial disadvantage faced by 3-proximal mutant alleles due to the binding partner effect is transient. However, the cumulative impact of this early disadvantage leads to a persistent suppression of the mutant allele relative to the wildtype, even by the time of pool stasis. We next discuss how the stalling effect[22, 30, 31, 21, 32, 24, 33, 23, 34, 35] can significantly prolong this transient phase, thereby amplifying the overall suppression of the mutant allele.
@@ III Stalling enhances collective suppression of mutants
Figure 6: Combination of collective effects and stalling suppresses mutant growth. (A) In duplexes where priming oligo has a correctly matched 3 end, extension proceeds normally and all such primers fully extend to form blunt ends. In duplexes where priming oligo has an incorrectly matched 3 end, extension past the mismatch is slowed by an additional time, $\tau_{stall}$. Consequently, only a fraction $1-e^{-\frac{\tau_{cycle}}{\tau_{stall}}}$ of mismatched oligos manage to extend within a thermal cycle of time $\tau_{cycle}$. See Supp. Sec. VII for further details. (B) (top) mut and wt amplification for three different values of $\frac{\tau_{stall}}{\tau_{cycle}}$ as a function of thermal cycle. (bottom) Mutant allele amplification measured after 10 thermal cycles as a function of $\frac{\tau_{stall}}{\tau_{cycle}}$.
The stalling effect describes the reduced rate of extension from a primer that ends in a mismatched base pair, compared to extension from a perfectly matched primer. Here, we investigate how stalling influences the propagation of mutant alleles within the VCG framework.
We consider annealed configurations where one of the oligos in the duplex is not correctly matched to its template at its 3 end. This can either happen when an oligo with a mutant allele at its 3 end is annealed to an oligo with a wt allele at that position, or it is an oligo with a wt allele at its 3 end annealed to an oligo with a mutant allele there. For such duplexes where one of the oligos has a mismatch at its 3 end, stalling delays extension by a time $\tau_{stall}$. Whether or not the duplex gets to extend therefore depends crucially on whether thermal cycling (with timescale $\tau_{cycle}$) drives the duplex apart before the stalling delay can be overcome. For primers with no 3 mismatch, there is no barrier to extension and such primers will fully extend during the current thermal cycle. To model the effect of stalling, we posit that primers with a mismatched 3 end get to extend through an exponential decay process with timescale $\tau_{stall}$. Hence by the end of a melt-anneal-extend cycle of length $\tau_{cycle}$, only a fraction $1-e^{-\frac{\tau_{cycle}}{\tau_{stall}}}$ of primers with a mismatched 3 end will have extended (Fig. 6A). Please see Supp. Sec. VII for further discussion and implementation details on stalled extension.
@@ IV Discussion
Our results show that collective replication in a virtual circular genome (VCG) architecture intrinsically suppresses the propagation of mutant alleles, even in the absence of high-fidelity copying or kinetic proofreading. Experiments, simulations, and a reduced theory show that suppression is strongest when mutations are 3-proximal and increases with the VCG’s ‘virtualness’ (redundancy of overlapping oligos).
@@
Intuitively, collective encoding of information builds a directional bias into replication. Alleles that match the consensus retain multiple routes to be replicated, whereas sequence changes (e.g., a mutant allele) prune those routes. We quantified these routes in terms of binding partners since an allele’s rate of being copied depends on how many annealed partners place their 3 end downstream of that allele so that extension traverses it. 3-proximal mutants have fewer such productive partners and thus are at a disadvantage compared to wildtype alleles that are coded at varying distances from the 3 end. Stalling at mismatched 3 ends prolongs this period of lowered replicative ability by delaying extension past the mismatch. Consequently, the combined effect of partner limitation and stalling leads to strong suppression that strengthens with virtualness and with increasing ratio $\tau_{stall}/\tau_{cycle}$ of stalling time to melt-anneal-extend cycle time. Unlike mispriming suppression in PCR, which relies on reduced binding affinity from mismatches, our effect persists even if mismatches have negligible impact on binding affinity.
@@
We focused on the maintenance of information in a structured pool, although prior works often focus on the emergence of catalytic cycles from random sequence pools[13, 14]. Prior classic theories of distributed replication (e.g., hypercycles[4]) posit cooperative autocatalytic networks that can maintain information, but they have been difficult to realize at the molecular level[36]. In contrast, our work begins from a concrete molecular specification — a set of overlapping oligos and their allowed anneal–extend interactions — which makes the system experimentally tractable and tunable. This molecular grounding, rather than an abstract mutually catalyzing graph, allowed us to quantify architecture-driven fidelity. In this aspect, our system shares features with molecularly-defined experimental models[37, 38, 39, 40, 41, 42].
We expect the core ecological mechanism revealed by our DNA model to generalize beyond the specific conditions studied here, though the quantitative outcomes will depend on the replication chemistry (e.g., in the context of RNA[10, 20]). Differences in persistence length, strand displacement rates, non-enzymatic extension kinetics will shift quantitative behavior. While we focused on primer extension, templated ligation is natural in distributed architectures [13, 14, 15, 43, 44, 12, 45, 46, 47] and enables growth from both 5 and 3 ends. The binding partner effect will have to be studied and quantified in the context of ligation to understand biases intrinsic to that mode of replication.
In addition to copying errors during primer extension, several other processes can introduce mutations that affect VCG stability. Mismatched annealing[48] can generate transient mispairs that alter extension outcomes, while sequence repeats[49] may create novel annealing registers, allowing oligomers to extend from unintended regions and producing chimeric products. Chemical lesions such as deamination[50] represent another important source of mutation, and these are not confined to the 3 terminus. Our systematic results on mutation position provide a framework to predict the fate of these diverse error sources within distributed architectures.
Our experiments emphasize primer extension in a fixed pool of oligomers that eventually become unproductive by forming long, unmeltable duplexes; sustained replication will require continual regeneration of short oligos (e.g., de novo synthesis from activated nucleotides[19], or cleaving oligos into short oligos[51]). Moving to driven continuous growth may strengthen mutant suppression, especially under strong stalling, because wildtype alleles can continue to amplify while 3-proximal mutants are effectively diluted out before they elongate sufficiently to act as a template[23, 24, 35, 22].
Despite these limitations and potential extensions, our work shows that genomic architecture can act as an intrinsic bias for the fate of stored information. That is, by filtering which novelties are propagated versus rejected, the architecture effectively tunes the error threshold without explicit proofreading. The Virtual Circular Genome model was proposed[19, 20] to explain how primitive replication systems could avoid issues associated with the replication of long linear genomes such as strand separation and replication of terminal bases. Surprisingly, we have now shown that it also mitigates the unrelated problem of high intrinsic error rates, and thereby provides a means of avoiding the Eigen error catastrophe[4] and thus allowing for the emergence of Darwinian evolution.
@@
Comparable architecture-level filters might be operating in extant biology, for example how RNA viruses with segmented genomes can undergo genetic reassortment[52]. Other non-standard genome architectures include those of multipartite bacteria (e.g., Borrelia burgdorferi[5, 53]), which partition essential genes across a chromosome plus several plasmids; ciliates package somatic information into vast numbers of gene-sized macronuclear chromosomes[7]; and trypanosomatids[8], which maintain kinetoplast DNA as a topologically linked network of maxi- and minicircles; and mitochondria and chloroplasts carry tens to thousands of genome copies per cell[6]. Even among close relatives, karyotypes differ sharply, e.g., Saccharomyces cerevisiae has 16 nuclear chromosomes whereas Schizosaccharomyces pombe has 3. In all these cases, information is distributed across many molecules with heterogeneous maintenance and inheritance, rather than on one continuous polymer. We lack a general framework for how such architectures influence mutation fate, dosage control, recombination, assortment noise, and long-term evolvability or why particular architectures arise and persist. While we do not address architectural features specific to these extant systems, our work highlights open questions about the advantages of distributing genomic information across many versus few molecules.